# MedDiff-FM: A Diffusion-based Foundation Model for Versatile Medical Image Applications

### v2

@@ 4.3 Evaluation of Downstream Tasks without Fine-tuning
Table 6: Anomaly detection performance on the liver CT anomaly detection from the BMAD benchmark.
@@
4.3.2 Image Denoising.
We evaluate the image denoising capabilities of MedDiff-FM on the Mayo 2016 dataset using the official 25% dose CT images. We utilize the off-the-shelf MedDiff-FM for image denoising as introduced in Section 3.4.1. The standard deviation of the additive noise $\sigma_{n}$ is estimated in the non-test cases and is around 0.15, while the regularization parameter $\lambda$ is set to 10. The results of the comparison methods in Table 5 are cited from [41]. Due to the different window width and level preferences of different methods, we report the PSNR and SSIM improvements of each method over its corresponding LDCT baseline under its respective window width and level, using the same test cases. Specifically, the LDCT baseline under the window of $[-1024,3072]$ HU yields a PSNR of 40.41 dB and an SSIM of 0.922, and the LDCT baseline under the window of $[-160,240]$ HU achieves a PSNR of 23.10 dB and an SSIM of 0.757. MedDiff-FM performs image denoising under the window of $[-160,240]$ HU. As shown in Table 5, MedDiff-FM even significantly outperforms supervised comparison methods in terms of SSIM, while maintaining competitive PSNR performance. Unlike all other comparison methods, MedDiff-FM is applied to image denoising without additional training on the Mayo 2016 dataset, thus reducing the consumption of spatio-temporal resources. The qualitative results in Fig. 6 intuitively demonstrate that MedDiff-FM achieves effective denoising and enhances image quality, highlighting its strong capabilities in image denoising.
@@
4.3.3 Anomaly Detection.
We utilize pre-trained MedDiff-FM to accomplish the anomaly detection task on the liver CT dataset from the BMAD [2] benchmark. Since the test images are 2D slices, totaling 1,493 slices, we stack them along the z-axis to form pseudo-3D volumes as inputs to MedDiff-FM during inference. As introduced in Section 3.4.2, MedDiff-FM adds noise to abnormal CT scans at a fixed time step to disrupt tumor structures. Because small time steps struggle to corrupt tumor regions [75], we diffuse the abnormal CT scans using a fixed time step of 950. The quantitative results of the baseline methods in Table 6 are taken from BMAD [2]. In accordance with [2], threshold-independent metrics are adopted for evaluation, with the Dice score being additionally used as a threshold-dependent reference metric. The quantitative results are summarized in Table 6. The pre-trained MedDiff-FM outperforms state-of-the-art anomaly detection methods by a large margin, especially in terms of Image AUROC, PRO, and Dice. Fig. 7 visualizes the anomaly maps, where MedDiff-FM accurately localizes liver tumors. These results highlight the strong potential of MedDiff-FM for anomaly detection in medical imaging.
Table 7: Super-resolution performance on the RPLHR-CT dataset.
@@
4.4 Evaluation of Downstream Tasks with Fine-tuning
Table 8: Quantitative comparison of lesion generation performance across four datasets with different kinds of lesions.
@@ 4.4 Evaluation of Downstream Tasks with Fine-tuning
4.4.1 Super-Resolution.
MedDiff-FM is fine-tuned for volumetric super-resolution using the training set from the RPLHR-CT dataset [78]. The fine-tuned MedDiff-FM is then evaluated on the test set of 100 cases from the same dataset. As demonstrated in Section 3.5.1, the standard deviation of the additive noise $\sigma_{n}$ is set to 1.0 as discussed in [78] and the regularization parameter $\lambda$ is set to 1. The $t_{start}$ starts from $T$, with 100 Neural Function Evaluations (NFEs) using uniform skipping. The scale factor $sf$ is 5 in depth dimension since the real-paired thick-CTs have a slice thickness of 5 mm, while the thin-CTs have a slice thickness of 1 mm.
The results of the comparison methods in Table 7 are reported as in [78]. As shown in Table 7, MedDiff-FM achieves competitive performance on the real-paired RPLHR-CT dataset for the volumetric super-resolution task. We take advantage of DiffPIR [86] to adapt MedDiff-FM to this inverse problem. DiffPIR requires a known degradation operator, i.e., a known downsampling operator. However, for real-paired thick-CTs and thin-CTs, we can only approximate the downsampling operator, which differs from the actual degradation process. As a result, the performance of MedDiff-FM is somewhat limited. Nevertheless, as visualization in Fig. 8, MedDiff-FM achieves promising super-resolution results that are close to ground-truth images.
4.4.2 Lesion Generation.
We fine-tune MedDiff-FM for lesion generation using four datasets containing various types of lesions: MSD-Lung, MSD-Liver, MED-LN, and ABD-LN. We compare the fine-tuned MedDiff-FM with Med-DDPM [16] and MedDiff-FM trained from scratch. For fair comparison, we train Med-DDPM conditioned on anatomical structure masks and lesion masks. The quantitative results on four lesion datasets are listed in Table 8. Taking into account the results across all generative metrics, the MedDiff-FM framework exhibits notable advantages over the Med-DDPM method, although the fine-tuned MedDiff-FM does not consistently achieve the best performance in LPIPS, FID, and MMD. It is important to note that these generative metrics assess the quality of synthetic lesion images from a holistic perspective. However, for lesion generation, it is more crucial to emphasize the quality of the lesions themselves. To provide a better assessment of synthetic lesion quality, we employ segmentation models trained on real lesion data to segment synthetic lesions and compute the Dice score. The Dice score results in the last column demonstrate that the fine-tuned MedDiff-FM significantly surpasses other methods in high-quality lesion generation.
@@
4.4.3 Lesion Inpainting.
To further evaluate the realism and quality of the synthetic lesions themselves, we apply the fine-tuned MedDiff-FM to the lesion inpainting task. DiffTumor [8] is designed for tumor synthesis on healthy abdominal organs. To ensure a fair comparison, we retrain the diffusion model of DiffTumor on each dataset while using its released autoencoder model pre-trained on large-scale abdominal data. Since image inpainting does not change the holistic structure of the original image, we consider it unnecessary to evaluate the quality of the holistic image using general generative metrics. We employ segmentation models trained on real lesions to segment inpainted lesions and compute the Dice score. The quantitative results on four lesion datasets are presented in Table 9, and the qualitative results are illustrated in Fig. 10. Given that the autoencoder is pre-trained on abdominal data and the lesion synthesis on the MED-LN dataset is particularly challenging, DiffTumor fails to generate mediastinal lymph nodes. In summary, both qualitative and quantitative results indicate that the synthetic lesions of MedDiff-FM are more consistent with the distribution of real lesions, which validates the effectiveness of MedDiff-FM.
Table 9: Quantitative comparison of lesion inpainting performance.
@@ 5 Discussion
6 Conclusion
In conclusion, this paper presents MedDiff-FM, a diffusion-based foundation model that deals with a wide range of medical image tasks. MedDiff-FM utilizes 3D CT images from diverse publicly available datasets and focuses on multiple anatomical regions, overcoming the limitations of previous work that were constrained by specific anatomical regions and particular tasks. MedDiff-FM is capable of handling multi-level integrated image processing with position embedding to build multi-level spatial relationships, and using anatomical structures and regions as conditions. The pre-trained diffusion foundation model can seamlessly perform tasks such as image denoising, anomaly detection, and image synthesis. Furthermore, MedDiff-FM deals with super-resolution, lesion generation, and lesion inpainting through rapid fine-tuning via ControlNet with task-specific conditions. The experimental results highlight the effectiveness of MedDiff-FM, making it a valuable tool for various medical image applications.

### v3

@@ IV-C2 Image Denoising
TABLE VI: Anomaly detection performance on the liver CT anomaly detection from the BMAD benchmark.
@@
IV-C3 Anomaly Detection
We utilize pre-trained MedDiff-FM to accomplish the anomaly detection task on the liver CT dataset from the BMAD [56] benchmark. The test images consist of 1,493 2D slices, which we stack along the z-axis to form pseudo-3D volumes as inputs to MedDiff-FM during inference. As introduced in Section III-D2, MedDiff-FM adds noise to abnormal CT scans at a fixed time step to disrupt tumor structures. Because small time steps struggle to corrupt tumor regions [17], we diffuse the abnormal CT scans using a fixed time step of 950. The quantitative results of the baseline methods in Table VI are taken from BMAD [56]. In accordance with [56], threshold-independent metrics are adopted for evaluation, with the Dice score being additionally used as a threshold-dependent reference metric. The quantitative results are summarized in Table VI. The pre-trained MedDiff-FM outperforms state-of-the-art anomaly detection methods by a large margin, especially in terms of Image AUROC, PRO, and Dice. Fig. 7 visualizes the anomaly maps, where MedDiff-FM accurately localizes liver tumors. These results highlight the strong potential of MedDiff-FM for anomaly detection in medical imaging.
@@ IV-C3 Anomaly Detection
IV-D Evaluation of Downstream Tasks with Fine-tuning
TABLE VII: Super-resolution performance on the RPLHR-CT dataset.
@@ IV-D Evaluation of Downstream Tasks with Fine-tuning
TABLE VIII: Quantitative comparison of lesion generation performance across four datasets with different kinds of lesions.
@@
IV-D1 Super-resolution
MedDiff-FM is fine-tuned for volumetric super-resolution using the training set from the RPLHR-CT dataset [57]. The fine-tuned MedDiff-FM is then evaluated on the test set of 100 cases from the same dataset. As demonstrated in Section III-E1, the standard deviation of the additive noise $\sigma_{n}$ is set to 1.0 as discussed in [57] and the regularization parameter $\lambda$ is set to 1. The $t_{start}$ starts from $T$, with 100 Neural Function Evaluations (NFEs) using uniform skipping. The scale factor $sf$ is 5 in depth dimension since the real-paired thick-CTs have a slice thickness of 5 mm, while the thin-CTs have a slice thickness of 1 mm.
The results of the comparison methods in Table VII are reported as in [57]. As shown in Table VII, MedDiff-FM achieves competitive performance on the real-paired RPLHR-CT dataset for the volumetric super-resolution task. We take advantage of DiffPIR [39] to adapt MedDiff-FM to this inverse problem. DiffPIR requires a known degradation operator, i.e., a known downsampling operator. However, for real-paired thick-CTs and thin-CTs, we can only approximate the downsampling operator, which differs from the actual degradation process. As a result, the performance of MedDiff-FM is somewhat limited. Nevertheless, as visualization in Fig. 8, MedDiff-FM achieves promising super-resolution results that are close to ground-truth images.
IV-D2 Lesion Generation
We fine-tune MedDiff-FM for lesion generation using four datasets containing various types of lesions: MSD-Lung, MSD-Liver, MED-LN, and ABD-LN. We compare the fine-tuned MedDiff-FM with Med-DDPM [33] and MedDiff-FM trained from scratch. For fair comparison, we train Med-DDPM conditioned on anatomical structure masks and lesion masks. The quantitative results on four lesion datasets are listed in Table VIII. Taking into account the results across all generative metrics, the MedDiff-FM framework exhibits notable advantages over the Med-DDPM method, although the fine-tuned MedDiff-FM does not consistently achieve the best performance in LPIPS, FID, and MMD. It is important to note that these generative metrics assess the quality of synthetic lesion images from a holistic perspective. However, for lesion generation, it is more crucial to emphasize the quality of the lesions themselves. To provide a better assessment of synthetic lesion quality, we employ segmentation models trained on real lesion data to segment synthetic lesions and compute the Dice score. The Dice score results in the last column demonstrate that the fine-tuned MedDiff-FM significantly surpasses other methods in high-quality lesion generation.
@@ IV-D2 Lesion Generation
TABLE IX: Quantitative comparison of lesion inpainting performance.
@@
IV-D3 Lesion Inpainting
To further evaluate the realism and quality of the synthetic lesions themselves, we apply the fine-tuned MedDiff-FM to the lesion inpainting task. DiffTumor [2] is designed for tumor synthesis on healthy abdominal organs. To ensure a fair comparison, we retrain the diffusion model of DiffTumor on each dataset while using its released autoencoder model pre-trained on large-scale abdominal data. Since image inpainting does not change the holistic structure of the original image, we consider it unnecessary to evaluate the quality of the holistic image using general generative metrics. We employ segmentation models trained on real lesions to segment inpainted lesions and compute the Dice score. The quantitative results on four lesion datasets are presented in Table IX, and the qualitative results are illustrated in Fig. 10. Given that the autoencoder is pre-trained on abdominal data and the lesion synthesis on the MED-LN dataset is particularly challenging, DiffTumor fails to generate mediastinal lymph nodes. In summary, both qualitative and quantitative results indicate that the synthetic lesions of MedDiff-FM are more consistent with the distribution of real lesions, which validates the effectiveness of MedDiff-FM.
@@ V Discussion
VI Conclusion
In conclusion, this paper presents MedDiff-FM, a diffusion-based foundation model that deals with a wide range of medical image tasks. MedDiff-FM utilizes 3D CT images from diverse publicly available datasets and focuses on multiple anatomical regions, overcoming the limitations of previous work that were constrained by specific anatomical regions and particular tasks. MedDiff-FM is capable of handling multi-level integrated image processing with position embedding to build multi-level spatial relationships, and using anatomical structures and regions as conditions. The pre-trained diffusion foundation model can seamlessly perform tasks such as image denoising, anomaly detection, and image synthesis. Furthermore, MedDiff-FM deals with super-resolution, lesion generation, and lesion inpainting through efficient fine-tuning via ControlNet with task-specific conditions. The experimental results highlight the effectiveness of MedDiff-FM, making it a valuable tool for various medical image applications.